b) If 15.0 grams of calcium hydroxide is combined with 75.0 mL of 0.500 M HCl, how many grams of calcium chloride would be formed?

1
Expert's answer
2022-03-24T11:20:02-0400

Balanced equation :


Ca(OH)2(aq) + 2HCl(aq) → CaCl2(aq) + 2H2O(l)


1 mol Ca(OH)2 react6s with 2 mol HCl


Molar mass Ca(OH)2 = 74.09 g/mol


Mol Ca(OH)2 in 15.0 g = 15.0 g / 74.09 g/mol = 0.202 mol


This will react with 0.202 mol * 2 = 0.404 mol HCl


Mol HCl in 75 mL of 0.500 M HCl solution:


mol HCl = 75 mL / 1000 mL/L * 0.5 mol /L = 0.03750 mol HCl


You do not have sufficient HCl to react with all the Ca(OH)2 The HCl is limiting





The simple answer is that no calcium hydroxide is formed


But If I assume that you really want to ask :


What mass of CaCl2 was produced


From the equation


0.03750 mol HCl will produce 0.03750 /2 = 0.01875 mol CaCl2


Molar mass CaCl2 = 40.0 + 35.5*2 = 111 g/mol


Mass of 0.01875 mol CaCl2 = 0.01875 mol * 111 g/mol = 2.08 g


Answer: 2.08 g CaCl2 will be produced